OpenStudy (adrynicoleb):

How do you prove that a point on a perpendicular bisector is equidistant (at equal distances) from the endpoints of the segment it intersects?

OpenStudy (anonymous):

 1+2=180 2=180-1-----(1) 2+3=180 put value of eq 1 in eq 2 180-1+3=180 1=3 proved

OpenStudy (adrynicoleb):

But what theorem or postulate is that? @harsh123

OpenStudy (anonymous):

this is #3

OpenStudy (e.mccormick):

[¬º-°]¬ Oh no! The ghost of posting past!

OpenStudy (adrynicoleb):

o_o That looks more like a zombie xD

OpenStudy (e.mccormick):

Yah... ghost is (((༼•̫͡•༽))) and I did not like it, so I used zombie.